# Investigating the Fermi-Hubbard model by the Tensor-Backflow method

### v3

@@ III.2 Comparisons with previous results
Structure factors determine periods of the stripe order in the ground state. The charge and spin structure factors are defined as: $C(\textbf{k})=\frac{1}{M^{2}}\sum_{ij}e^{i\textbf{k}\cdot(\textbf{r}_{i}-\textbf{r}_{j})}n_{i}n_{j}$, $S(\textbf{k})=\frac{1}{M^{2}}\sum_{ij}e^{i\textbf{k}\cdot(\textbf{r}_{i}-\textbf{r}_{j})}s_{i}^{z}s_{j}^{z}$, where $n_{i}$ and $s_{i}^{z}$ are defined in Fig.(3). $C(\textbf{k})$ is set to zero when $k_{x}=k_{y}$=0. All results of structure factors are obtained by pp=1 wave-functions. Fig.(4) depicts structure factors for the filling of nn=0.875 under various $U$ and boundary conditions, on lattices of $16\times 16$ and $8\times 32$. As depicted by Fig.(4)(a)(b), charge structure factor peaks at $\pi/4$ and spin structure factor peaks at $7\pi/8$, corresponding to periods of 8 and 16 for CDW and SDW, respectively. Fig.(4)(c)(d) depicts results of a direct optimization on the $8\times 32$ lattice for nn=0.875 and UU=8 under PBC. A vertical stripe order is obtained, meanwhile structure factors peaks at same locations as on the $16\times 16$ lattice. On the $16\times 16$ lattice, the lattice’s side length of 16 corresponds to a whole period of the SDW, thus peaks on the $16\times 16$ lattice are higher than those on the $8\times 32$ lattice.
Tab.(5) depicts additional energy results for nn=0.875 and UU=8 under PBC for both $t^{\prime}$=0 and $t^{\prime}$=-0.2. All results in the table are obtained by direct optimizations. For $t^{\prime}$=0 on the $8\times 16$ lattice, the vertical stripe order is obtained with same periods as those on the $16\times 16$ lattice.
In particular, energies for $t^{\prime}$=-0.2 are depicted in Tab.(5). The $t^{\prime}$=-0.2 is important for understanding mechanisms of superconductivity Hubbard_SC1 . From energy comparisons in Tab.(5), we find that NNN backflow (QQ=9) achieves slightly lower energy than NN backlfow (QQ=5) for the case of $t^{\prime}$=-0.2 on the $8\times 16$ lattice. Because of the non-zero NNN hopping strengths, NNN backflow are considered for all cases of $t^{\prime}$=-0.2.
@@
Because cases of $t^{\prime}$=-0.2 are more challenging than those of $t^{\prime}$=0, optimizations are easily to get trapped on local minima, especially on large lattices. We show that choosing a proper initial UHF is very helpful to avoid the local minima for the final converged energy Hubbard_NN4 . In Tab.6, we compare energies converged by Tensor-Backflow with initial UHFs from different strengths of $U$ on the $12\times 12$ lattice. For cases of UU=8, $U_{\mathrm{UHF}}$=4 leads to the best converged energy of Tensor-Backflow. Ground state patterns on the $12\times 12$ lattice for various $U_{\mathrm{UHF}}$ are depicted in Appendix.D.
On the $8\times 16$ lattice, the ground state pattern for $t^{\prime}$=-0.2 is depicted in Fig.(5), with definitions of average spin and density per site from Fig.(3). For the case on the $8\times 16$ lattice, the initial UHF is insufficiently optimized under $U_{\mathrm{UHF}}=U$. From Fig.(5), a width-4 stripe is achieved, yielding to the result on the same lattice from NQS Hubbard_NN3 . In the optimization of the NQS, energies were compared between vertical and horizontal stripes and the horizontal stripe achieved lower energy Hubbard_NN3 . In this work, the horizontal stripe is achieved naturally after optimizations without pinning-fields.
In Fig.(6), we compare spin structure factors between $t^{\prime}$=0 and -0.2. To reval detailed differences, we increase the number of k-points from $17^{2}$ to $65^{2}$. Thus, there are totally 33 k-points in the x-axis. From the figure, for $t^{\prime}$=0 peaks are sharply distributed and centered on $7\pi/8$. However, the peak for $t^{\prime}$=-0.2 is wider comparing to that for $t^{\prime}$=0. The wider peak reveals competitions between stripe orders of different periods in the ground state for $t^{\prime}$=-0.2, thus the optimization is easily to get stuck on a local minima. The high optimization difficulty of $t^{\prime}$=-0.2 is depicted by Tab.(6) on the $12\times 12$ lattice. Therefore, comparisons of various $U_{\mathrm{UHF}}$ are necessary for $t^{\prime}$=-0.2 on the $16\times 16$ lattice in near future.
@@
For other fillings such nn=0.8 and nn=0.9375, energy results are presented on Tab.(7). As depicted by cases of nn=0.8, UU=8 on the $10\times 10$ lattice, the improvement of all-site backflow is moderate comparing to NN backflow. Thus, for a Hamiltonian with NN hoppings, the major state representation is achieved by the NN backflow.
Fig.(7) depicts charge and spin structure factors for nn=0.8 on the $10\times 20$ lattice and for nn=0.9375 on the $8\times 32$ lattice. Based on the phase diagram via AFQMC AFQMC_phase_diagram , the critical interaction strength $U_{c}$ between un-ordered and ordered ground states increase with the doping level $h$. Here, for nn=0.8, the spin structure factor peaks at $4\pi/5$ for UU=8, meanwhile there are no obvious peaks for UU=4. Comparing charge structure factors between UU=4 and UU=8, there is the tendency of the obvious peak in the charge structure factor when further increasing $U$. For nn=0.9375, the spin structure factor peaks at $15\pi/16$, meanwhile there are no obvious peaks in the charge structure factor. Results of nn=0.8 and nn=0.9375 in Fig.(7) yield to the phase diagram via AFQMC AFQMC_phase_diagram ; AFQMC_n0.9375 .
@@ IV Discussions and conclusions
In this work, we benchmarked the Tensor-Backflow method along with a Lanczos optimization for various fillings $n$, interaction strengths $U$, NNN hopping strengths $t^{\prime}$, lattice sizes and boundary conditions. Comparing to current state-of-the-art methods such as fPEPS Hubbard_fPEPS , NQS Hubbard_NN3 ; Hubbard_NN4 , achieved energies are competitive or even lower. In addition to energy results, we investigate ground state properties through patterns of $s_{i}^{z}$ and $n_{i}$ as well as charge and spin structure factors. When $t^{\prime}$=0, obtained state properties are consistent with other state-of-the-art methods such as fPEPS Hubbard_fPEPS and NQS Hubbard_NN3 for nn=0.875 as well as AFQMC for nn=0.8 AFQMC_phase_diagram and nn=0.9375 AFQMC_n0.9375 . For $t^{\prime}$=-0.2 on the $8\times 16$ lattice, the width-4 horizontal stripe is achieved, matching the result from NQS Hubbard_NN3 . On the $12\times 12$ lattice, achieved energy is lower than the state-of-the-art NQS result Hubbard_NN3 , meanwhile a width-4 stripe is obtained based on $n_{i}$.
@@
From our results, although a Lanczos optimization significantly improves the energy, the property of the state is determined by the pp=0 wave-function. Namely, a Lanczos optimization can’t override the property of the original wave-function, as patterns of $s_{i}^{z}$ and $n_{i}$ are similar between pp=0 and pp=1 wave-functions. For example, Fig.(8) depicts density-density correlations from pp=0 and pp=1 wave-functions for nn=0.875 and $U=8$ on the $8\times 32$ lattice under PBC. From the figure, the CDW with the period of 8 is already in the pp=0 wave-function and there are no significant differences for pp=1 results. Therefore, differences between correlations from pp=0 and pp=1 wave-functions are negligible.
Pair correlations determine the existence of superconductivitiy tJ_fPEPS ; pair_pair_correlation_1 ; pair_pair_correlation_2 . In this work, based on comparisons between pp=0 and pp=1 wave-functions, to avoid high calculation complexity, we evaluate $d$-wave pair correlations based on pp=0 wave-functions, with pair correlations defined in the literature tJ_fPEPS . For the case of nn=0.875 and UU=8 on the $16\times 16$ lattice under PBC, when $t^{\prime}$=0 and $t^{\prime}$=-0.2, pair correlations are in the magnitude of $\mathcal{O}(10^{-5})$ when the distance between two pairs is $L/2$, where $L$ is the side length of the square lattice. Because optimizations are easily to get trapped on local minima for cases of $t^{\prime}$=-0.2, more careful and intensive investigations are necessary in the future work.
Based on results in this work, the Tensor-Backflow is an efficient state representation for the lattice Fermi-Hubbard model. Furthermore, there are several ways to improve the energy precision. For example, considering more backflow terms comparing to hopping terms in the Hamiltonian, comparing energies under various initializations, comparing ground state patterns with and without applying pinning-fields, applying more advanced optimization methods such as the stochastic reconfiguration, etc. Currently, the wave-function is not symmetry enforced, thus translational invariance is not preserved during optimizations. For PBC, the energy precision can be further improved by enforcing translational invariance on the wave-function. The Tensor-Backflow has no restrictions on lattice shapes and boundary conditions, thus it is potentially an efficient and universal method for solving lattice fermion systems.
Figure 9: Comparisons of ground state patterns between two optimization strategies in Sec.III.1. The case is nn=0.875, UU=8 and $t^{\prime}$=0 on the $4\times 16$ lattie under OBC, with NN backflow in wave-functions. Left-sided figures denote the average spin per site: $s_{i}^{z}=(n_{i}^{\uparrow}-n_{i}^{\downarrow})/2$, and right-sided figures denote the average density per site: $n_{i}=n_{i}^{\uparrow}+n_{i}^{\downarrow}$. (a)(b) and (c)(d) are achieved by optimization strategies 1 and 2, respectively.
@@ Appendix A Comparing ground state patterns of two optimization strategies
Fig.(9) compares ground state patterns between two optimization strategies in Sec.III.1, and results are from pp=0 wave-functions. The case is nn=0.875, UU=8 on the $4\times 16$ lattice under OBC. From the figure, both optimizations strategies achieve similar ground state patterns. Meanwhile, ground state patterns in Fig.(9) yield to those from state-of-the-art NQS results Hubbard_NN3 .
@@ Appendix B Comparisons between initial UHFs and subsequent Tensor-Backflow wave-functions
We show that Tensor-Backflow achieves an essentially different state from the initial UHF by structure factors $C(\mathbf{k})$ and $S(\mathbf{k})$. For nn=0.875 on the $16\times 16$ lattice under PBC, cases of UU=8,10,12 are depicted in Fig.(10), Fig.(11) and Fig.(12), respectively. Based on definitions in Sec.III.3, correlations in both $C(\mathbf{k})$ and $S(\mathbf{k})$ are classical, thus they are calculated through averages on MC samples. Here, each structure factor is obtained by 46080 MC samples, results of UHF and Tensor-Backflow are from pp=0 and pp=1 wave-functions, respectively.
In our calculations, each MC sample is flattened from two-dimensional to one-dimensional in the same way and each struture factor is calcualted by the same procedure from MC samples. Comparing (a)(b) to (c)(d) in Fig.(10), (11) and (12), properties of the converged wave-function is not simply inherited from the initial UHF. Furthermore, the Tensor-Backflow can override properties of the initial UHF. In each case, there is $U_{\mathrm{UHF}}=U$, and initial UHFs are not fully optimized in all three cases.
@@ Appendix C Energy convergence of transferring from UU=8
Fig.(13) depicts energy convergences when transferring the converged wave-function for UU=8 and PBC to other cases. Because of the good initial state, required optimization step number is much less than that from a direct optimization. For cases denoted by (a), (b) and (c) in Fig.(13), energies achieved by direct optimizations are -0.6771, -0.6264 and -0.6771, respectively. Meanwhile, energies achieved by transferring from UU=8 are -0.6787, -0.6272 and -0.6786, respectively.

### v4

@@ III.2 Comparisons with previous results
Structure factors provide insight into the periods of stripe order in the ground state. The charge and spin structure factors are defined as: $C(\textbf{k})=\frac{1}{M^{2}}\sum_{ij}e^{i\textbf{k}\cdot(\textbf{r}_{i}-\textbf{r}_{j})}n_{i}n_{j}$, $S(\textbf{k})=\frac{1}{M^{2}}\sum_{ij}e^{i\textbf{k}\cdot(\textbf{r}_{i}-\textbf{r}_{j})}s_{i}^{z}s_{j}^{z}$, where $n_{i}$ and $s_{i}^{z}$ are defined in Fig.(3). Note that $C(\mathbf{k})$ is set to zero when $k_{x}=k_{y}=0$. All results for structure factors are computed using the pp=1 wave-functions. Fig.(4) shows the structure factors for a filling of nn=0.875 at various interaction strengths $U$ and boundary conditions, on lattices of $16\times 16$ and $8\times 32$. As seen in Fig.(4)(a) and (b), the charge structure factor peaks at $\pi/4$, and the spin structure factor peaks at $7\pi/8$, corresponding to periods of 8 and 16 for charge-density-wave (CDW) and spin-density-wave (SDW) orders, respectively. Fig.(4)(c) and (d) present results from a direct optimization on the $8\times 32$ lattice for nn=0.875 and UU=8 under PBC. A vertical stripe order is observed, with structure factor peaks at the same locations as those on the $16\times 16$ lattice. On the $16\times 16$ lattice, the side length of 16 corresponds to a full period of the SDW, which results in higher peak values compared to those on the $8\times 32$ lattice.
Tab.(5) presents additional energy results for nn=0.875 and UU=8 under PBC for both $t^{\prime}$=0 and $t^{\prime}$=-0.2. All values in the table are obtained through direct optimizations. For $t^{\prime}$=0 on the $8\times 16$ lattice, a vertical stripe order emerges, with the same periodicity observed on the $16\times 16$ lattice.
Energy results for $t^{\prime}$=-0.2 are shown in Tab.(5), which plays a crucial role in understanding the mechanisms of superconductivity Hubbard_SC1 . From the comparisons presented in the table, we observe that NNN backflow (QQ=9) yields slightly lower energies than NN backflow (QQ=5) for the $t^{\prime}$=-0.2 case on the $8\times 16$ lattice. Due to the non-zero NNN hopping strengths, NNN backflow is considered for all calculations involving $t^{\prime}$=-0.2.
@@
The cases with $t^{\prime}$=-0.2 are more challenging than those with $t^{\prime}$=0, as optimizations are more easy to getting trapped in local minima, particularly for larger lattices. We demonstrate that selecting an appropriate initial UHF is crucial for avoiding such local minima and achieving the best final energy Hubbard_NN4 . Tab.(6) compares the energies converged by Tensor-Backflow with initial UHFs corresponding to different values of $U$ on the $12\times 12$ lattice. For UU=8, an initial UHF with $U_{\mathrm{UHF}}$=4 leads to the best converged energy. Ground state patterns for various $U_{\mathrm{UHF}}$ values on the $12\times 12$ lattice are presented in App.( D).
On the $8\times 16$ lattice, the ground state pattern for $t^{\prime}$=-0.2 is shown in Fig.(5), with the definitions of the average spin and charge density per site as given in Fig.(3). For this case, the initial UHF is insufficiently optimized when $U_{\mathrm{UHF}}$=U. As shown in Fig.(5), a width-4 stripe pattern emerges, consistent with the result obtained on the same lattice by NQS Hubbard_NN3 . In the NQS optimization, energy comparisons between vertical and horizontal stripe patterns were made, with the horizontal stripe yielding a lower energy Hubbard_NN3 . In contrast, in this work, the horizontal stripe naturally emerges after optimization, without the need for pinning-fields.
In Fig.(6), we compare the spin structure factors for $t^{\prime}$=0 and $t^{\prime}$=-0.2. To highlight the differences more clearly, we increase the number of $k$-points from $17^{2}$ to $65^{2}$, resulting in a total of 33 $k$-points along the $x$-axis. From the figure, we observe that for $t^{\prime}$=0, the peaks are sharply defined and centered at $7\pi/8$. However, for $t^{\prime}$=-0.2, the peak is broader compared to the $t^{\prime}$=0 case. This broader peak indicates competition between stripe orders with different periods in the ground state for $t^{\prime}$=-0.2, which makes the optimization more easy to getting trapped in local minima. The increased optimization difficulty for $t^{\prime}$=-0.2 is further illustrated in Tab.(6) for the $12\times 12$ lattice. Therefore, comparisons across various values of $U_{\mathrm{UHF}}$ will be necessary for $t^{\prime}$=-0.2 on the $16\times 16$ lattice in future studies.
@@
For other fillings, such as nn=0.8 and nn=0.9375, energy results are provided in Tab.(7). In the case of nn=0.8 and UU=8 on the $10\times 10$ lattice, the improvement from all-site backflow over NN backflow is relatively moderate. Therefore, for a Hamiltonian with NN hoppings, the primary contribution to the state representation comes from the NN backflow.
Fig.(7) shows the charge and spin structure factors for nn=0.8 on the $10\times 20$ lattice and for nn=0.9375 on the $8\times 32$ lattice. According to the phase diagram from AFQMC AFQMC_phase_diagram , the critical interaction strength $U_{c}$ that separates the disordered and ordered ground states increases with the doping level $h$. For nn=0.8, the spin structure factor for UU=8 peaks at $4\pi/5$, whereas for UU=4, no obvious peak is observed. In the charge structure factor, an increasing peak appears as $U$ increases, especially at UU=8. For nn=0.9375, the spin structure factor peaks at $15\pi/16$, but no clear peak is seen in the charge structure factor. The results for both nn=0.8 and nn=0.9375 in Fig.(7) are consistent with the phase diagram from AFQMC AFQMC_phase_diagram ; AFQMC_n0.9375 .
@@ IV Discussions and conclusions
In this work, we benchmark the Tensor-Backflow method, followed by a Lanczos step, across a range of fillings $n$, interaction strengths $U$, NNN hopping strengths $t^{\prime}$, lattice sizes, and boundary conditions. Compared to state-of-the-art methods such as fPEPS Hubbard_fPEPS and NQS Hubbard_NN3 ; Hubbard_NN4 , the energies achieved by Tensor-Backflow are competitive or even lower. In addition to energy results, we investigate the ground state properties, including the spin ($s_{i}^{z}$) and density ($n_{i}$) patterns, as well as the charge and spin structure factors. For $t^{\prime}$=0, the state properties obtained by Tensor-Backflow are consistent with those from fPEPS Hubbard_fPEPS and NQS Hubbard_NN3 for nn=0.875, and with AFQMC for nn=0.8 AFQMC_phase_diagram and nn=0.9375 AFQMC_n0.9375 . For $t^{\prime}$=-0.2 on the $8\times 16$ lattice, a width-4 horizontal stripe is obtained, consistent with the result from NQS Hubbard_NN3 . On the $12\times 12$ lattice, the achieved energy is lower than the best-known NQS result Hubbard_NN3 , with a width-4 stripe identified based on the charge density $n_{i}$.
Our results show that, although a Lanczos step significantly improves the energy, the properties of the state are determined primarily by the pp=0 wave-function. Specifically, a Lanczos step does not change the fundamental properties of the original wave-function, as the patterns of $s_{i}^{z}$ and $n_{i}$ remain similar between the pp=0 and pp=1 wave-functions. For instance, Fig.(8) shows the density-density correlations from both the pp=0 and pp=1 wave-functions for nn=0.875 and $U=8$ on the $8\times 32$ lattice under PBC. As seen in the figure, the CDW with a period of 8 is already present in the pp=0 wave-function, and no significant differences are observed in the correlations for the pp=1 wave-function. Therefore, the differences between the two-body correlations from the pp=0 and pp=1 wave-functions are negligible.
@@
Figure 9: Comparisons of ground state patterns between two optimization strategies in Sec. (III.1). The case is nn=0.875, UU=8 and $t^{\prime}$=0 on the $4\times 16$ lattie under OBC, with NN backflow in wave-functions. Left-sided figures denote the average spin per site: $s_{i}^{z}$, and right-sided figures denote the average density per site: $n_{i}$. (a)(b) and (c)(d) are achieved by optimization strategies 1 and 2, respectively.
Pair correlations are essential for determining the presence of superconductivity tJ_fPEPS ; pair_pair_correlation_1 ; pair_pair_correlation_2 . In this work, to mitigate the computational complexity, we evaluate $d$-wave pair correlations using the pp=0 wave-functions, based on the definition of pair correlations in the literature tJ_fPEPS . For the case of nn=0.875 and UU=8 on the $16\times 16$ lattice under PBC, we find that pair correlations are on the order of $\mathcal{O}(10^{-5})$ when the distance between pairs is $L/2$, where $L$ is the side length of the square lattice, for both $t^{\prime}$=0 and $t^{\prime}$=-0.2. Since optimizations are easy to getting trapped in local minima, particularly for cases with $t^{\prime}$=-0.2, the future work includes comparing energies with various stripe orders by applying pinning-fields and calculating pair correlations using pp=1 wave-functions.
Based on the results presented in this work, Tensor-Backflow provides an efficient state representation for the lattice Fermi-Hubbard model. There are several methods for improving energy precision. For instance, compare energies for different initializations or investigate ground state patterns both with and without the application of pinning fields. At present, the wave-function is not symmetry-enforced, meaning translational invariance is not preserved during optimization. For systems with PBC, energy precision could be further enhanced by enforcing translational invariance in the wave-function. Additionally, the Tensor-Backflow method is flexible with respect to lattice shapes and boundary conditions, making it a potentially efficient and universal approach for solving lattice fermion systems.
@@ Appendix A Comparing ground state patterns of two optimization strategies
Fig.(9) compares the ground state patterns obtained from the two optimization strategies described in Sec. (III.1), using $p$ = 0 wave-functions. The case corresponds to nn=0.875, UU=8 on the $4\times 16$ lattice under OBC. From the figure, both optimization strategies lead to similar ground state patterns. Moreover, the patterns in Fig.(9) are in agreement with those from state-of-the-art NQS results Hubbard_NN3 .
@@ Appendix B Comparisons between initial UHFs and subsequent Tensor-Backflow wave-functions
We demonstrate that the Tensor-Backflow method produces a state that is fundamentally different from the initial UHF, as revealed by the charge and spin structure factors $C(\mathbf{k})$ and $S(\mathbf{k})$. For nn=0.875 on the $16\times 16$ lattice under PBC, we show results for UU=8, 10, and 12 in Fig.(10), Fig.(11), and Fig.(12), respectively. As described in Sec. (III.3), the correlations in both $C(\mathbf{k})$ and $S(\mathbf{k})$ are classical in nature, and are therefore computed by averaging over MC samples. Specifically, each structure factor is calculated using 46,080 MC samples, with the results for the UHF and Tensor-Backflow states derived from the pp=0 and pp=1 wave-functions, respectively.
In our calculations, each MC sample is flattened from two-dimensional to one-dimensional using the same procedure, and each structure factor is computed consistently from the MC samples. By comparing panels (a) and (b) with (c) and (d) in Fig.(10), Fig.(11), and Fig.(12), we observe that the properties of the converged wave-function are not simply inherited from the initial UHF. Moreover, the Tensor-Backflow method allows the wave-function to override the properties of the initial UHF. In all cases, we set $U_{\mathrm{UHF}}=U$, and the initial UHFs are not fully optimized in any of the three scenarios.
@@ Appendix C Energy convergence of transferring from UU=8
Fig.(13) shows the energy convergence when transferring the converged wave-function from UU=8 and PBC to other cases. Due to the quality of the initial state, the number of optimization steps required for transfer is significantly smaller compared to direct optimization. For the cases labeled (a), (b), and (c) in Fig.(13), the energies obtained by direct optimization are -0.6771, -0.6264, and -0.6771, respectively. In contrast, the energies achieved by transferring the wave-function from UU=8 are -0.6787, -0.6272, and -0.6786, respectively.